# Environmentally Sustainable Software Design and Development: A Systematic Literature Review

### v1

@@ 6. Related Work
Several secondary studies survey the state of the art for sustainable software. A first example is given by Ormazabal et al (Ormazabal et al., 2014), who survey the software tools that can be used to assess a product’s lifecycle and analyse its carbon footprint. Sinha and Chandel (Sinha and Chandel, 2014) and Sudarmadi and Garniwa (Sudarmadi and Garniwa, 2023) review the use of software tools for planning of hybrid renewable energy systems. Popescu et al. (Popescu et al., 2022) survey the existing approaches for building digital twins of manufacturing production systems by incorporating environmental sustainability aspects. The above listed studies differ from ours in their objective, as they discuss how to achieve environmental sustainability with software. We instead focus on the solutions enhancing the environmental sustainability of software itself.
@@
Other noteworthy secondary studies are the literature reviews by Salam and Khan (Salam and Khan, 2016) and Sambhanthan and Potdar (Sambhanthan and Potdar, 2016). Salam and Khan (Salam and Khan, 2016) classify and analyse the success factors for vendors of green software. Sambhanthan and Potdar (Sambhanthan and Potdar, 2016) instead focus on the existing frameworks for assessing a business’ sustainability, therein included ICT sustainability. Therefore, the reviews by Salam and Khan (Salam and Khan, 2016) and Sambhanthan and Potdar (Sambhanthan and Potdar, 2016) differ from ours in the objective, as they focus on the business-related aspects of software’s sustainability. We instead aim at classifying the technical solutions for designing and developing environmentally sustainable software.
@@
Andrikopoulos et al. (Andrikopoulos et al., 2022) provide a systematic mapping study aimed at analysing which architecting activities are affected/affect which of the four different dimensions of sustainability (i.e., technical, economic, environmental, and social). Ahmadisakha and Andrikopoulos (Ahmadisakha and Andrikopoulos, 2024) deepen into the topic, specifically focusing Cloud-based software architecture, still analysing how architectural decisions for Cloud software are affected/affect the four different dimensions of sustainability. Therefore, the scope and goal of both the above secondary studies differs from ours, as they aim at eliciting which of the four different dimensions of sustainability impact on/are impacted by which architecting phase. We instead focus on the 5W1H of designing and developing environmentally sustainable software. Additionally, Ahmadisakha and Andrikopoulos (Ahmadisakha and Andrikopoulos, 2024) focus on Cloud-based software, while we rather consider multiple different application domains.
@@
Other secondary studies focus on enhancing the environmental sustainability of software in specific target application domains. For example, Azimzadeh and Tabrizi (Azimzadeh and Tabrizi, 2015) and Khan et al. (Khan et al., 2020) provide two different reviews on software techniques for managing green Cloud data centres. Gaglianese et al. (Gaglianese et al., 2023b) survey the existing approaches for the environmentally sustainable orchestration of containerised software applications over Cloud-Edge infrastructures. Castaño et al. (Castaño et al., 2023) analyse and classify the carbon footprint of training publicly available ML models. Verdecchia et al. (Verdecchia et al., 2023) review the state of the art on green AI, by distinguishing the proposals supporting the training and operation of AI models. López-Pérez et al. (López-Pérez et al., 2022) survey the existing techniques for operating 5G networks in an energy-efficient manner. Thus, all the above-mentioned secondary studies focus on a specific application domain and on different lifecycle stages than ours, as we rather consider designing and developing environmentally sustainable software by covering – and classifying – multiple different application domains.
@@
The SLRs by Bozzelli et al. (Bozzelli et al., 2014), Calero et al. (Calero et al., 2013), Koziolek(Koziolek, 2011), and Venters et al. (Venters et al., 2018, 2023) are a step closer to ours, as they all pertain to the broader scopus of environmental sustainability in software engineering. More precisely, Bozzelli et al. (Bozzelli et al., 2014) and Koziolek (Koziolek, 2011) classify the existing metrics for estimating the greenness of software. Calero et al. (Calero et al., 2013) elicit the software quality aspects related to environmental sustainability. Venters et al. (Venters et al., 2018, 2023) go a step further, by also trying to identify which definitions of software sustainability are available, as well as how software sustainability impacts on architectural decisions, goals, and education. Thus, even if slightly overlapping, the objectives of the above mentioned SLRs differ from ours, as we rather aim at answering to the 5W1H questions for designing and developing environmentally sustainable software (see Section 2.1).
@@
Similar arguments apply to the SLR by Penzenstadler et al. (Penzenstadler et al., 2012), who started with similar aims than ours, to then realise that – at the time when the study was conducted – only a few studies tackled environmental sustainability in software engineering. The SLR by Penzenstadler was then redirected to identify and classify the broader scope of relating environmental sustainability to software, one of these being how it relates with software engineering. Therefore, the SLR by Penzenstadler et al. (Penzenstadler et al., 2012) differs from ours in the ultimate contribution, as – after more than 10 years – the literature on sustainability is more mature, as we showed by focusing on the narrower scope of designing and developing environmentally sustainable software.
@@
In summary, to the best of our knowledge, ours is the first SLR on the state of the art on designing and developing environmentally sustainable software. It is the first doing it considering not only the what of sustainable software design and development, but rather aiming at answering to the all 5W1H questions related to designing and developing environmentally sustainable software (see Section 2.1). This enabled us, e.g., to first reason on the stakeholders for environmentally sustainable software design and development, to span over multiple application domains, and to first distinguish whether environmental sustainability is dealt with by only reducing energy consumption or by also considering the carbon footprint of software.
@@ 7. Conclusions and Outlook
We presented the results of a SLR on the design and development of environmentally sustainable software. More precisely, after classifying and recapping the existing solutions for the tasks, we analysed who are the targeted stakeholders, what methods are proposed to design and develop environmentally sustainable software, why, where and when the proposed solutions are applied (in terms of application domains and lifecycle stages, respectively), and how the selected studies have been presented to the scientific community.
@@
The results of our SLR can help researchers and practitioners identify and adopt existing guidelines and reference models for designing and developing environmentally sustainable software, as well as to select applicable solutions for their software projects. Our analysis also highlights a growing interest in this topic since 2020, providing a basis for establishing future research directions. To facilitate this, we outline seven research directions derived from our analysis, each mapped to the highlights in Section 4 that motivate these directions.
@@
Relieving the burden from software developers (–). The existing solutions mainly aim at supporting software developers, and this somehow puts the burden of designing and developing environmentally sustainable software on software developers themselves. There is a need for solutions that could support and involve all the potential stakeholders in the process. In particular, we should enable software architects and sustainability engineers to get more involved in the design of environmentally sustainable software, taking responsibility for sustainability-oriented design choices, thus complementing the work and efforts carried out by software developers.
@@
Standardising environmental sustainability for software (). Most of the surveyed proposals are reference models, witnessing the need for clarifying the notion of environmental sustainability for software, and how this can be used to support the design and development of more sustainable software. However, despite sharing some base concepts, the currently available reference models differ one from another, calling for standardisation efforts. The availability of a standardised reference model for software environmental sustainability would enable all stakeholders to rely on the same understanding of sustainability. It would also lay the foundations for devising full-fledged solutions to support the design and development of environmentally sustainable software.
@@
Finding a trade-off between applicability and sustainability enhancements (). Most of the existing solutions for designing and developing environmentally sustainable software adopt a black-box approach. This is mainly motivated by applicability reasons, as black-box approaches can work even with software whose sources are not available, e.g., for privacy or legacy reasons. At the same time, while surveying the state of the art, we observed some potentials of white-box analyses. For example, Le Goaër et al (Goaër, 2021, 2023; Goaër and Hertout, 2022) highlighted the potential of defining, detecting, and resolving so-called green code smells, namely possible symptoms of bad coding decisions that may worsen the environmental impact of software. Mehra et al. (Mehra et al., 2023) instead showcased how static analysis enabled them to successfully resolve code inefficiencies in an industrial use case. Therefore, the potentials of white-box approaches deserve further investigation, in an effort to find a suitable trade-off between the applicability of a solution and its effectiveness in supporting the design and development of environmentally sustainable software.
@@
Raising the overall TRL (). We observed that, overall, the TRL of the state of the art on environmentally sustainable software design and development is relatively low (with only one out of 65 selected studies reaching a TRL 5). To favour the adoption of existing – and newly proposed – solutions by the software industry, the overall TRL should be raised. In particular, the existing proposals should be prototyped and assessed over industry-relevant scenarios, to then release more mature tooling to the software industry. To this end, there is also a need for benchmarks, industry-relevant use cases, and realistic case studies that could be used to thoroughly validate existing/newly proposed solutions for designing and developing environmentally sustainable software.
@@
Increasing carbon-awareness (). Most of the existing solutions consider energy consumption as the main discriminant to enhance the environmental sustainability of software. This is motivated by the shared assumption that, by directly reducing the energy consumption of software, its carbon footprint gets indirectly reduced as well. However, energy mixes play a crucial role to relate energy consumption with carbon footprint, as different energy mixes come with different associated carbon intensities (Aiello et al., 2024). This may impair the effectiveness of directly considering energy consumption to indirectly reduce carbon emission, since e.g., when software is powered with clean and renewable energy, reducing its energy consumption does not significantly change its carbon footprint (Forti et al., 2024). This calls for further research along the line of directly considering energy mixes and carbon emissions of software, rather than only focussing on reducing its carbon footprint by limiting its energy consumption.
@@
Adopting existing solutions in application domains (,–). Most of the existing solutions are high-level guidelines, reference models, and measurement solutions for generic software. Being them generic, their concrete adoption in specific application domains may require considerable efforts, e.g., to adapt them to work with the peculiar aspects that mostly impact on the carbon footprint on application domain-specific software (like the training phase for ML models, or the PUE for cloud data centers). This calls for identifying and considering such peculiar aspects, to define/improve domain-specific guidelines and reference models, as well as to adapt existing techniques to work application domain-specific software. This would enable the effective use and assessment of existing solutions available in multiple different application domains, thus possibly paving the way towards their adoption. Additionally, narrowing the scope of to a specific application domains showed potentials to not only reduce, but also minimise, the environmental impact of the software under design/development.
@@
Relating design and development with all lifecycle stages (–). Even if focussing on the design and development of environmentally sustainable software, this SLR highlighted how the existing solutions span across all the stages of the software lifecycle. This is motivated by the fact that sustainability guidelines and reference models are often proposed for the full software lifecycle. Going beyond this, we observed that the effectiveness of the design and development stages in achieving environmental sustainability impacts on/is impacted by the other lifecycle stages as well. This calls for supporting the software design and development stages by considering, keeping track, and assessing their impacts/relations with the whole software lifecycle.
@@ Acknowledgements.
- Ahmed (2020) Shakeel Ahmed. 2020. Environmental sustainability coding techniques for cloud computing. International Journal of Advanced Computer Science and Applications 11, 10 (2020), 231 – 237. https://doi.org/10.14569/IJACSA.2020.0111031
@@
- Andrikopoulos et al. (2022) Vasilios Andrikopoulos, Rares-Dorian Boza, Carlos Perales, and Patricia Lago. 2022. Sustainability in Software Architecture: A Systematic Mapping Study. In 48th Euromicro Conference on Software Engineering and Advanced Applications, SEAA 2022, Maspalomas, Gran Canaria, Spain, 31 August - 2 September 2022, Gustavo Marrero Callicó, Regina Hebig, and Andreas Wortmann (Eds.). IEEE, 426–433. https://doi.org/10.1109/SEAA56994.2022.00073
@@
- Azimzadeh and Tabrizi (2015) Aryan Azimzadeh and Nasseh Tabrizi. 2015. A Taxonomy and Survey of Green Data Centers. In 2015 International Conference on Computational Science and Computational Intelligence (CSCI). 128–131. https://doi.org/10.1109/CSCI.2015.70
@@
- Beghoura et al. (2014) Mohamed Amine Beghoura, Abdelhak Boubetra, and Abdallah Boukerram. 2014. Green Applications Awareness: NonLinear Energy Consumption Model for Green Evaluation. In Eighth International Conference on Next Generation Mobile Apps, Services and Technologies, NGMAST 2014, Oxford, United Kingdom, September 10-12, 2014. IEEE, 48–53. https://doi.org/10.1109/NGMAST.2014.33
@@
- Brownlee et al. (2021) Alexander E. I. Brownlee, Jason Adair, Saemundur O. Haraldsson, and John Jabbo. 2021. Exploring the Accuracy - Energy Trade-off in Machine Learning. In 10th IEEE/ACM International Workshop on Genetic Improvement, GI@ICSE 2021, Madrid, Spain, May 30, 2021. IEEE, 11–18. https://doi.org/10.1109/GI52543.2021.00011
@@
- Calero et al. (2013) Coral Calero, Manuel F. Bertoa, and María Ángeles Moraga. 2013. A systematic literature review for software sustainability measures. In 2nd International Workshop on Green and Sustainable Software, GREENS 2013, San Francisco, CA, USA, May 20, 2013, Patricia Lago, Niklaus Meyer, Maurizio Morisio, Hausi A. Müller, and Giuseppe Scanniello (Eds.). IEEE Computer Society, 46–53. https://doi.org/10.1109/GREENS.2013.6606421
@@
- Condori-Fernández and Lago (2018) Nelly Condori-Fernández and Patricia Lago. 2018. The Influence of Green Strategies Design onto Quality Requirements Prioritization. In Requirements Engineering: Foundation for Software Quality - 24th International Working Conference, REFSQ 2018, Utrecht, The Netherlands, March 19-22, 2018, Proceedings (Lecture Notes in Computer Science, Vol. 10753), Erik Kamsties, Jennifer Horkoff, and Fabiano Dalpiaz (Eds.). Springer, 189–205. https://doi.org/10.1007/978-3-319-77243-1_12
- Danushi et al. (2024) Ornela Danushi, Stefano Forti, and Jacopo Soldani. 2024. Environmentally Sustainable Software Design and Development: Replication Package. Zenodo, https://doi.org/10.5281/zenodo.12807332.
@@
- Dick et al. (2013) Markus Dick, Jakob Drangmeister, Eva Kern, and Stefan Naumann. 2013. Green software engineering with agile methods. In 2nd International Workshop on Green and Sustainable Software, GREENS 2013, San Francisco, CA, USA, May 20, 2013, Patricia Lago, Niklaus Meyer, Maurizio Morisio, Hausi A. Müller, and Giuseppe Scanniello (Eds.). IEEE Computer Society, 78–85. https://doi.org/10.1109/GREENS.2013.6606425
- Dick et al. (2010) Markus Dick, Stefan Naumann, and Norbert Kuhn. 2010. A Model and Selected Instances of Green and Sustainable Software. In What Kind of Information Society? Governance, Virtuality, Surveillance, Sustainability, Resilience - 9th IFIP TC 9 International Conference, HCC9 2010 and 1st IFIP TC 11 International Conference, CIP 2010, Held as Part of WCC 2010, Brisbane, Australia, September 20-23, 2010. Proceedings (IFIP Advances in Information and Communication Technology, Vol. 328), Jacques Berleur, Magda David Hercheui, and Lorenz M. Hilty (Eds.). Springer, 248–259. https://doi.org/10.1007/978-3-642-15479-9_24
@@
- Fu et al. (2021) Andre Fu, Mahdi S. Hosseini, and Konstantinos N. Plataniotis. 2021. Reconsidering CO2 Emissions From Computer Vision. In IEEE Conference on Computer Vision and Pattern Recognition Workshops, CVPR Workshops 2021, virtual, June 19-25, 2021. Computer Vision Foundation / IEEE, 2311–2317. https://doi.org/10.1109/CVPRW53098.2021.00262

### v2

@@ 6. Related Work
Several secondary studies survey the state of the art for sustainable software. A first example is given by Ormazabal et al (Ormazabal et al., 2014), who survey the software tools that can be used to assess a product’s lifecycle and analyse its carbon footprint. Sinha and Chandel (Sinha and Chandel, 2014) and Sudarmadi and Garniwa (Sudarmadi and Garniwa, 2023) review the use of software tools for planning of hybrid renewable energy systems. Popescu et al. (Popescu et al., 2022) survey the existing approaches for building digital twins of manufacturing production systems by incorporating environmental sustainability aspects. The above listed studies differ from ours in their objective, as they discuss how to achieve environmental sustainability with software. We instead focus on the solutions improving the environmental sustainability of software itself.
@@
Other noteworthy secondary studies are the literature reviews by Salam and Khan (Salam and Khan, 2016) and Sambhanthan and Potdar (Sambhanthan and Potdar, 2016). Salam and Khan (Salam and Khan, 2016) classify and analyse the success factors for vendors of green software. Sambhanthan and Potdar (Sambhanthan and Potdar, 2016) instead focus on the existing frameworks for assessing a business’ sustainability, therein included ICT sustainability. Therefore, the reviews by Salam and Khan (Salam and Khan, 2016) and Sambhanthan and Potdar (Sambhanthan and Potdar, 2016) differ from ours in the objective, as they focus on the business-related aspects of software’s sustainability. We instead aim at classifying the technical solutions for designing and developing carbon-efficient software.
@@
Andrikopoulos et al. (Andrikopoulos et al., 2022) provide a systematic mapping study aimed at analysing which architecting activities are affected/affect which of the four different dimensions of sustainability (i.e., technical, economic, environmental, and social). Ahmadisakha and Andrikopoulos (Ahmadisakha and Andrikopoulos, 2024) deepen into the topic, focussing specifically on Cloud-based software architecture, still analysing how architectural decisions for Cloud software are affected/affect the four different dimensions of sustainability. Therefore, the scope and goal of both the above secondary studies differ from ours, as they aim at eliciting which of the four different dimensions of sustainability impact on/are impacted by which architecting phase. We instead focus on the 5W1H of designing and developing carbon-efficient software. Additionally, Ahmadisakha and Andrikopoulos (Ahmadisakha and Andrikopoulos, 2024) focus on Cloud-based software, while we rather consider multiple different application domains.
@@
Other secondary studies focus on improving the environmental sustainability of software in specific target application domains. For example, Azimzadeh and Tabrizi (Azimzadeh and Tabrizi, 2015) and Khan et al. (Khan et al., 2020) provide two different reviews on software techniques for managing green Cloud data centres. Gaglianese et al. (Gaglianese et al., 2023b) survey the existing approaches for the environmentally sustainable orchestration of containerised software applications over Cloud-Edge infrastructures. Castaño et al. (Castaño et al., 2023) analyse and classify the carbon footprint of training publicly available ML models. Verdecchia et al. (Verdecchia et al., 2023) review the state of the art on green AI, by distinguishing the proposals supporting the training and operation of AI models. López-Pérez et al. (López-Pérez et al., 2022) survey the existing techniques for operating 5G networks in an energy-efficient manner. Thus, all the above-mentioned secondary studies focus on a specific application domain and on different lifecycle stages than ours, as we rather consider designing and developing carbon-efficient software by covering – and classifying – multiple different application domains.
@@
The SLRs by Bozzelli et al. (Bozzelli et al., 2014), Calero et al. (Calero et al., 2013), Koziolek(Koziolek, 2011), and Venters et al. (Venters et al., 2018, 2023) are a step closer to ours, as they all pertain to the broader scopus of environmental sustainability in software engineering. More precisely, Bozzelli et al. (Bozzelli et al., 2014) and Koziolek (Koziolek, 2011) classify the existing metrics for estimating the greenness of software. Calero et al. (Calero et al., 2013) elicit the software quality aspects related to environmental sustainability. Venters et al. (Venters et al., 2018, 2023) go a step further, by also identifying which definitions of software sustainability are available, as well as how software sustainability impacts on architectural decisions, goals, and education. Thus, even if slightly overlapping, the objectives of the above mentioned SLRs differ from ours, as we rather aim at answering to the 5W1H questions for designing and developing carbon-efficient software (see Section 2.1).
@@
Similar arguments apply to the SLR by Penzenstadler et al. (Penzenstadler et al., 2012), who started with similar aims than ours, to then realise that – at the time when the study was conducted – only a few studies tackled environmental sustainability in software engineering. The SLR by Penzenstadler was then redirected to identify and classify the broader scope of relating environmental sustainability to software, one of these being how it relates with software engineering. Therefore, the SLR by Penzenstadler et al. (Penzenstadler et al., 2012) differs from ours in the ultimate contribution, as – after more than 10 years – the literature on sustainability is more mature, as we showed by focusing on the narrower scope of designing and developing carbon-efficient software.
@@
In summary, to the best of our knowledge, ours is the first SLR on the state of the art on designing and developing carbon-efficient software. It is the first doing it considering not only the what of sustainable software design and development, but rather aiming at answering to the all 5W1H questions related to designing and developing carbon-efficient software (see Section 2.1). This enabled us, e.g., to first reason on the stakeholders for carbon-efficient software design and development, to span over multiple application domains, and to first distinguish whether environmental sustainability is dealt with by only reducing energy consumption or by also considering the carbon footprint of software.
@@ 7. Conclusions and Outlook
We presented the results of a SLR on the design and development of carbon-efficient software. After classifying and recapping the existing solutions for the tasks, we analysed who are the targeted stakeholders, what methods allow to design and develop carbon-efficient software, why, where and when the proposed solutions are applied (in terms of application domains and lifecycle stages, respectively), and how the selected studies have been presented to the scientific community.
@@
The results of our SLR can help researchers and practitioners identify and adopt existing guidelines and reference models for designing and developing carbon-efficient software, as well as to select applicable solutions for their software projects. Our analysis also highlights a growing interest in this topic since 2020, providing a basis for establishing future research directions. To facilitate this, we outline seven research directions derived from our analysis, each mapped to the highlights in Section 4 that motivate these directions.
@@
Relieving the burden from software developers (–). The existing solutions mainly aim at supporting software developers, and this somehow puts the burden of designing and developing carbon-efficient software on software developers themselves. There is a need for solutions to support and involve all the potential stakeholders in the process. In particular, we should enable software architects and sustainability engineers to get more involved in the design of carbon-efficient software, taking responsibility for sustainability-oriented design choices, thus complementing the work and efforts carried out by software developers.
@@
Standardising carbon efficiency for software (). Most of the surveyed proposals are reference models, witnessing the need for clarifying the notion of carbon efficiency for software, also outlining how this can be used to support the design and development of more sustainable software. However, despite sharing some base concepts, the currently available reference models differ one from another, calling for standardisation efforts. The availability of a standardised reference model for software carbon efficiency would enable all stakeholders to rely on the same understanding of carbon efficiency. It would also lay the foundations for devising full-fledged solutions to support the design and development of carbon-efficient software.
@@
Finding a trade-off between applicability and carbon efficiency (). Most of the existing solutions for designing and developing carbon-efficient software adopt a black-box approach. This is mainly motivated by applicability reasons, as black-box approaches can work even with software whose sources are not available, e.g., for privacy or legacy reasons. At the same time, while surveying the state of the art, we observed some potentials of white-box analyses. For example, Le Goaër et al (Goaër, 2021, 2023; Goaër and Hertout, 2022) highlighted the potential of defining, detecting, and resolving so-called green code smells, namely possible symptoms of bad coding decisions that may worsen the environmental impact of software. Mehra et al. (Mehra et al., 2023) instead showcased how static analysis enabled them to successfully resolve code inefficiencies in an industrial use case. Therefore, the potentials of white-box approaches deserve further investigation, in an effort to find a suitable trade-off between the applicability of a solution and its effectiveness in supporting the design and development of carbon-efficient software.
@@
Raising the overall TRL (). We observed that, overall, the TRL of the state of the art on carbon-efficient software design and development is relatively low (with only one out of 65 selected studies reaching a TRL 5). To favour the adoption of existing – and newly proposed – solutions by the software industry, the overall TRL should be increased. In particular, the existing proposals should be prototyped and assessed over industry-relevant scenarios, to then release more mature tooling to the software industry. To this end, there is also a need for benchmarks, industry-relevant use cases, and realistic case studies that could be used to thoroughly validate existing/newly proposed solutions for designing and developing carbon-efficient software.
@@
Increasing carbon-awareness (). Most of the existing solutions consider energy consumption as the main discriminant to improve the environmental sustainability of software. This is motivated by the shared assumption that, by directly reducing the energy consumption of software, its carbon footprint gets indirectly reduced as well. However, energy mixes play a crucial role to relate energy consumption with carbon footprint, as different energy mixes come with different associated carbon intensities (Aiello et al., 2024). This may impair the effectiveness of directly considering energy consumption to indirectly reduce carbon emission, since e.g., when software is powered with clean and renewable energy, reducing its energy consumption does not significantly change its carbon footprint (Forti et al., 2024). This calls for further research along the line of directly considering energy mixes and carbon emissions of software, rather than only focussing on reducing its carbon footprint by limiting its energy consumption.
@@
Adopting existing solutions in application domains (,–). Most of the existing solutions are high-level guidelines, reference models, and measurement solutions for generic software. Being them generic, their concrete adoption in specific application domains may require considerable efforts, e.g., to adapt them to work with the peculiar aspects that mostly impact on the carbon footprint on application domain-specific software (like the training phase for ML models, or the PUE for Cloud data centers). This calls for identifying and considering such peculiar aspects, to define/improve domain-specific guidelines and reference models, as well as to adapt existing techniques to work with application domain-specific software. This would enable the effective use and assessment of existing solutions available in multiple different application domains, thus possibly paving the way towards their adoption. Additionally, narrowing the scope to specific application domains showed the potential to not only reduce, but also minimise, the carbon efficiency of the software under design/development.
@@
Relating design and development with all lifecycle stages (–). Even if focussing on the design and development of carbon-efficient software, this SLR highlighted how the existing solutions span across all the stages of the software lifecycle. This is motivated by the fact that sustainability guidelines and reference models are often proposed for the full software lifecycle. Going beyond this, we observed the effectiveness of the design and development stages in achieving carbon-efficient impacts on/is impacted by the other lifecycle stages as well. This calls for supporting the software design and development stages by considering, keeping track of, and assessing their impacts/relations with the whole software lifecycle.
Additionally, although this survey focuses specifically on the design and development stages, the outcomes discussed in Section 4.5 highlight that other lifecycle stages might contribute to the carbon-efficiency of software artifacts, e.g., requirements engineering might set carbon-efficiency goals, testing might help validating carbon-efficiency, and operation might further reduce the carbon footprint of running software. While the carbon-efficient operation of software has already been analysed, e.g., in Gaglianese et al. (Gaglianese et al., 2023b), future secondary studies could complement the results of this work by exploring how the other lifecycle stages can further contribute to the carbon-efficiency of software artifacts. Such secondary studies are left in the scope of future work.
@@ Acknowledgements.
- Ahmed (2020) Shakeel Ahmed. 2020. Environmental sustainability coding techniques for cloud computing. Int. Journal of Advanced Computer Science and Applications 11, 10 (2020), 231 – 237. https://doi.org/10.14569/IJACSA.2020.0111031
@@
- Andrikopoulos et al. (2022) Vasilios Andrikopoulos, Rares-Dorian Boza, Carlos Perales, and Patricia Lago. 2022. Sustainability in Software Architecture: A Systematic Mapping Study. In 48th Euromicro Conference on Software Engineering and Advanced Applications, SEAA 2022, Maspalomas, Gran Canaria, Spain, 31 August - 2 September 2022, Gustavo Marrero Callicó et al. (Eds.). IEEE, 426–433. https://doi.org/10.1109/SEAA56994.2022.00073
@@
- Azimzadeh and Tabrizi (2015) Aryan Azimzadeh and Nasseh Tabrizi. 2015. A Taxonomy and Survey of Green Data Centers. In 2015 International Conf. on Computational Science and Computational Intelligence (CSCI). 128–131. https://doi.org/10.1109/CSCI.2015.70
@@
- Beghoura et al. (2014) Mohamed Amine Beghoura, Abdelhak Boubetra, and Abdallah Boukerram. 2014. Green Applications Awareness: NonLinear Energy Consumption Model for Green Evaluation. In Eighth Int. Conf. on Next Generation Mobile Apps, Services and Technologies, NGMAST 2014, Oxford, United Kingdom, September 10-12, 2014. IEEE, 48–53. https://doi.org/10.1109/NGMAST.2014.33
@@
- Brownlee et al. (2021) Alexander E. I. Brownlee, Jason Adair, Saemundur O. Haraldsson, and John Jabbo. 2021. Exploring the Accuracy - Energy Trade-off in Machine Learning. In 10th IEEE/ACM Int. Workshop on Genetic Improvement, GI@ICSE 2021, Madrid, Spain, May 30, 2021. IEEE, 11–18. https://doi.org/10.1109/GI52543.2021.00011
@@
- Calero et al. (2013) Coral Calero, Manuel F. Bertoa, and María Ángeles Moraga. 2013. A systematic literature review for software sustainability measures. In 2nd International Workshop on Green and Sustainable Software, GREENS 2013, San Francisco, CA, USA, May 20, 2013, Patricia Lago et al. (Eds.). IEEE Computer Society, 46–53. https://doi.org/10.1109/GREENS.2013.6606421
@@
- Condori-Fernández and Lago (2018) Nelly Condori-Fernández and Patricia Lago. 2018. The Influence of Green Strategies Design onto Quality Requirements Prioritization. In Requirements Engineering: Foundation for Software Quality - 24th Int. Working Conf., REFSQ 2018, Utrecht, The Netherlands, March 19-22, 2018, Proceedings (Lecture Notes in Computer Science, Vol. 10753), Erik Kamsties et al. (Eds.). Springer, 189–205. https://doi.org/10.1007/978-3-319-77243-1_12
- Danushi et al. (2025) Ornela Danushi, Stefano Forti, and Jacopo Soldani. 2025. Carbon-Efficient Software Design and Development: Replication Package. Zenodo, https://doi.org/10.5281/zenodo.14917055.
@@
- Dick et al. (2013) Markus Dick, Jakob Drangmeister, Eva Kern, and Stefan Naumann. 2013. Green software engineering with agile methods. In 2nd Int. Workshop on Green and Sustainable Software, GREENS 2013, San Francisco, CA, USA, May 20, 2013, Patricia Lago et al. (Eds.). IEEE Computer Society, 78–85. https://doi.org/10.1109/GREENS.2013.6606425
- Dick et al. (2010) Markus Dick, Stefan Naumann, and Norbert Kuhn. 2010. A Model and Selected Instances of Green and Sustainable Software. In What Kind of Information Society? Governance, Virtuality, Surveillance, Sustainability, Resilience - 9th IFIP TC 9 Int. Conf., HCC9 2010 and 1st IFIP TC 11 Int. Conf., CIP 2010, Held as Part of WCC 2010, Brisbane, Australia, September 20-23, 2010. Proceedings (IFIP Advances in Information and Communication Technology, Vol. 328), Jacques Berleur et al. (Eds.). Springer, 248–259. https://doi.org/10.1007/978-3-642-15479-9_24
@@
- Fu et al. (2021) Andre Fu, Mahdi S. Hosseini, and Konstantinos N. Plataniotis. 2021. Reconsidering CO2 Emissions From Computer Vision. In IEEE Conf. on Computer Vision and Pattern Recognition Workshops, CVPR Workshops 2021, virtual, June 19-25, 2021. Computer Vision Foundation / IEEE, 2311–2317. https://doi.org/10.1109/CVPRW53098.2021.00262